40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 185 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  7786   Tue Dec 4 20:38:51 2012 jamieOmnistructureComputersnew (beta) version of nds2 installed on control room machines

I've installed the new nds2 packages on the control room machines.

These new packages include some new and improved interfaces for python, matlab, and octave that were not previously available. See the documentation in:

  /usr/share/doc/nds2-client-doc/html/index.html

for details on how to use them.  They all work something like:

  conn = nds2.connection('fb', 8088)
  chans = conn.findChannels()
  buffers = conn.fetch(t1, t2, {c1,...})
  data = buffers(1).getData()

NOTE: the new interface for python is distinct from the one provided by pynds.  The old pynds interface should continue to work, though.

To use the new matlab interface, you have to first issue the following command:

   javaaddpath('/usr/lib/java')

I'll try to figure out a way to have that included automatically.

The old Matlab mex functions (NDS*_GetData, NDS*_GetChannel, etc.) are now provided by a new and improved package.  Those should now work "out of the box".

  7785   Tue Dec 4 20:13:55 2012 KojiUpdateSUSTTs are ready

Please leave here what was the instruction by Bran and Suresh so that the other people can redo it sometime later!

  7784   Tue Dec 4 18:53:50 2012 AyakaUpdateAlignmentBS chamber

BS chamber seemed to be kicked again around 10:00 am today.

I moved PZT mainly in YAW and locked both arms. I adjusted the beam to be almost on the center of both ETM by sights.

BSchamb1204.pdf

  7783   Tue Dec 4 18:06:35 2012 DenUpdateSUSTTs are ready

 Using instructions from Bram and Suresh, I was able to plug in connectors to BOSEMs. Today I've tested electronics, everything works good. Jamie made an medm screen and channels for TTs. Sliders for pitch and yaw go from -100 to 100 counts. Calibration to angle is 1e-5 rad / count.

TTs are in the clean room waiting for installation.

IMG_0105.JPG    IMG_0108.JPG

  7782   Tue Dec 4 11:30:36 2012 ZachUpdatePSLPMC calibration for MC_F calibration

In order to have less unknown, you can calibrate the PMC PZT separately. Lock the PMC and take a transfer function from either the NPRO PZT input or the FSS AOM VCO input to the PMC control signal. The VCO is better, since the calibration should be much better known, but I am not sure what the current setup of the 40m PSL is, so I don't know if the FSS is normally locked.

Since you know the NPRO PZT or VCO actuation coefficients, you can assume the PMC loop (where the OL gain is high enough) is correcting for the frequency fluctuations. So, simply multiply the known coefficient by the transfer function to get the PMC PZT gain.

Then, you can re-do your PMC PZT sweep measurement and be confident of the calibration. The FSR must be right, so you can get the finesse with confidence.

Quote:

Hmm... Does anyone find falses in my measurement?
If not, the finesse can be 4 times smaller than the value which was 5 years ago?

 

  7781   Tue Dec 4 10:51:10 2012 SteveUpdateSUSPRM damping restored

PRM oplev gains set to zero from PIT 0.15 and YAW -0.3 and damping restored

  7780   Tue Dec 4 10:44:36 2012 SteveUpdatePEMshort power outage

 

CALIFORNIA INSTITUTE OF TECHNOLOGY

                 FACILITIES MANAGEMENT

 

**PLEASE POST**

 

 

Building:         Campus

 

Date:             Thursday October 04,2012

 

This morning at 2:17 a.m. much of the City of Pasadena including our Campus experienced a electric power sag of short duration, approximately 1/10 of a second. The cause was a fault on one of Pasadena’s 17KV circuits. Some sensitive equipment have been impacted.

                 

Contact:          Mike Anchondo x-4999

  7779   Tue Dec 4 01:43:37 2012 ranaUpdatePSLPMC calibration for MC_F calibration

  If you can't find the PMC sidebands in the transmission, its because the SNR is too small.

It may be a better idea to look at the PMC error signal, since the DC signal there is suppressed by the demodulation.

  7778   Mon Dec 3 17:04:12 2012 AyakaUpdatePSLPMC calibration for MC_F calibration

In order to calibrate MC_F signal, I need to know the calibration value from thorlab's PZT driver to laser frequency.
The calibration value should be ~ 15MHz/V (the PZT driver has 15 gain, and the laser has the calibration value of ~ 1MHz/V according to the laser spec sheet), but I want to confirm it.
This can be measured by sweeping the input voltage of the PZT driver and see the transmission signal from unlocked PMC.
 

1. Response of PMC transmission when the signal is inputted to laser PZT

I inputted 0.2 Hz triangular wave with 5Vamp and 2.5V offset into the PZT driver and see the transmission signal from PMC. After the PZT driver and before the laser, there is an analog low pass filter but its cut off frequency is 1 Hz so I did not take it into consideration. 
DSC_4955.JPGsweep_PZTdriver.png(TEK00000.CSV, TEK00001.CSV in the zip file)

I could not the side-band resonances. I guess it was because the generated signal is not big enough (but still the maximum range of the signal generator.)
Therefore, in order to calibrate the input voltage to the frequency, I need to know finesse or FWHM frequency.
 

2. Responce of PMC transmission when the voltage of PZT on the PMC is swept

In order to measure the finesse and FWHM frequency, I also swept the PMC PZT voltage with the DC offset slider at the FSS.adl and tried to measure the finesse of PMC. (reference: elog #904)
PMC_PZT_FSR.pdftrans_fit.png(PMC-PZTcal_121203.xml in the zip file)

The result of fitting:

V_FSR (the PZT voltage difference between the 2 resonances) ~ 63 +/- 7 V (= 731MHz (given))
V_FWHM (the PZT voltage to sweep FWHM) ~  0.32 +/- 0.04 V (~ 3.7 MHz)
Finesse ~ 200 +/- 30

However, this finesse value is much smaller than the value on the Wiki, 800. (Manasa showed me.)
V_FSR is comparable to the result Rana got at the referenced elog. But I am not sure about the V_FWHM because it is hard to figure out how large the PZT voltage changed from the template file (PMC-PZTcal.xml).
Are those mode wrong? But if so, where is the correct mode resonances? I think they should be visible...
 

3. Calibration value 

When I know the FWHM frequency, I can calibrate the input on the PZT driver into laser frequency.

The results are:

if I take the finesse of 800 and FSR of 731 MHz (the values on the Wiki): ~5.0 MHz/V
if I take the finesse of 200 and FSR of 731 MHz (the measured value): ~20.0 MHz/V

Actually, the measured value is closer to the value calculated from the spec sheet.

Hmm... Does anyone find falses in my measurement?
If not, the finesse can be 4 times smaller than the value which was 5 years ago?

Attachment 5: PMC-PZTcal_121203.zip
  7777   Mon Dec 3 16:37:09 2012 SteveUpdatePEMEM 172 microphones ordered

Quote:

I've put EM 172 microphones inside Steve's isolation box to measure their noise. I've attached mics to each other and aligned them using the tape.

At low frequencies (below 1 Hz) the noise is limited by ADC as there is a 10 Hz high-pass filter inside mic readout box.

ADC noise is measured by splitting the signal from 1 mic into 2 ADC channels.

em172.png

 BT EM172 microphones  are ordered.

  7776   Mon Dec 3 07:40:00 2012 SteveHowToAlignmenthow was the BS chamber misalinged

Quote:

 

 Whoever was working around the BS chamber at 11 AM on Friday should admit it now and take the punishment.

For those of you who like to do work on the interferometer without reporting it in the elog because you think that what you did doesn't affect anything, this is your example of how our time can be wasted by such laziness.

 I'm taking full responsibility for this action and I told them after lunch Friday.

 

HOW NOT TO:

The BS isolation stack  supported by two beam tubes and they can pivot around the pivot point.

Attachment 1: BSsupport.jpg
BSsupport.jpg
  7775   Sun Dec 2 00:37:49 2012 DenUpdateSUSTT cable problem

 This week I've got all TT stuff baked and today was testing eddy current damping and electronics.

In the beginning everything was good: ring magnets fit mirror holder holes and their interaction with actuation magnets is strong enough to keep damping magnets in the wholes. I've put the frame horizontally and kicked it, magnets were still in the whole. Brackets also fit to the TT frame.  

DSC_4945.JPG     DSC_4946.JPG

I've tested eddy current dumping during ring down measurements, it was strong enough.

DSC_4947.JPG    DSC_4948.JPG

Then I started to test electronics. I've provided signal to TT1 channels and could see it in the clean room. But then things went terrible. I just could not connect TT cables to OSEMS, there is not enough space in the OSEM for the connector to plug in.

DSC_4949.JPG     DSC_4952.JPG

Connector should be machines to be more narrow. There is actually no reason for a connector to have this shape. I think it was designed to fit perfectly the OSEM frame but turned out to be ~0.5 mm wider then it should be.

  7774   Sat Dec 1 16:58:14 2012 ManasaUpdateWIKI-40M UpdateOptical tables

I have updated the wiki with the layout of the out-of-vac optical tables: Updated optical tables

I used the new camera to take pictures.

Lesson learnt after the update:

To use the new canon to take better pictures of optics tables; set the camera to manual mode; no flash and iso at around 800 or higher if you can hold the camera still for that long. The autofocus works beautifully...so you will not need any minor tweaking of lens to take pictures. 

  7773   Sat Dec 1 13:56:38 2012 ranaUpdateAlignmentBS chamber

 

 Whoever was working around the BS chamber at 11 AM on Friday should admit it now and take the punishment.

For those of you who like to do work on the interferometer without reporting it in the elog because you think that what you did doesn't affect anything, this is your example of how our time can be wasted by such laziness.

  7772   Sat Dec 1 00:24:37 2012 Den, AyakaUpdateAlignmentBS chamber

Today at 11:13 AM the stack of invacuum BS table was kicked and IFO misaligned. We adjusted PZT2 voltage by ~20 V in yaw such that IPPOS was restored. Then we could lock arms.

push.png

  7771   Sat Dec 1 00:13:16 2012 DenUpdateAdaptive FilteringARMS and MC

Quote:

 

 I actuate on ETMY for YARM and ETMX for XARM. For now I did adaptive filtering for both arms at the same time. I used the same parameters for xarm as for yarm.

I've notched 16 Hz resonance because it has high Q and I need to think more how to subtract it using FIR filter or apply IIR.

I'll try MC stabilazation method.

 Adaptive filtering was applied to MC and X,Y arms at the same time. I used a very aggressive (8 order) butterworth filter at 6 Hz as an AI filter for MC not to inject noise to ARMS as was done before

Mu for MC was 0.2, downsample = 16, delay = 1. I was able to subtract 1 Hz. Stack subraction is not that good as for arms but this is because I used only one seismometer for MC that is under the BS. I might install accelerometers under MC2.

EDIT, JCD, 18Feb2013:  Den remembers using mu for the arms in the range of 0.01 to 0.1, although using 0.1 will give extra noise.  He said he usually starts with something small, then ramps it up to 0.04, and after it has converged brings it back down to 0.01.

Attachment 1: arms_mcl_oaf.pdf
arms_mcl_oaf.pdf
  7770   Fri Nov 30 23:10:36 2012 CharlesUpdateElectronicsVertex Illuminators

 3 of the 4 remote controlled illuminators at the vertex are installed and can now be turned on via sitemap. There are a total of 15 controls for "Illum", but only the 3 labeled with MC, BS-PRM and ITMY-SRM are functional.

  7769   Fri Nov 30 22:11:50 2012 DenUpdateAdaptive FilteringARMS

Quote:

This is interesting. I suppose you are acting on the ETMY.
Can you construct the compensation filter with actuation on the MC length?
Also can you see how the X arm is stabilized?

This may stabilize or even unstabilize the MC length, but we don't care as the MC locking is easy.

If we can help to reduce the arm motion with the MCL feedforward trained with an arm sometime before,
this means the lock acquisition will become easier. And this may still be compatible with the ALS.

Why did you notched out the 16Hz peak? It is the dominant component for the RMS and we want to eliminate it.

 I actuate on ETMY for YARM and ETMX for XARM. For now I did adaptive filtering for both arms at the same time. I used the same parameters for xarm as for yarm.

I've notched 16 Hz resonance because it has high Q and I need to think more how to subtract it using FIR filter or apply IIR.

I'll try MC stabilazation method.

Attachment 1: arms_oaf.pdf
arms_oaf.pdf
  7768   Fri Nov 30 14:21:18 2012 ranaHowToComputer Scripts / ProgramsThe mystery of PDFs and you. As deep as the mystery of Rasputin.

This is how to post PDF:

From DTT, print the plot as a postscript file.

Then use ps2pdf to make a archival PDF version (the flag is the key!). Example:

ps2pdf -dPDFX /home/controls/Desktop/darm.ps

Attachment 1: darm.pdf
darm.pdf
  7767   Fri Nov 30 11:49:24 2012 KojiUpdateAdaptive FilteringYARM

This is interesting. I suppose you are acting on the ETMY.
Can you construct the compensation filter with actuation on the MC length?
Also can you see how the X arm is stabilized?

This may stabilize or even unstabilize the MC length, but we don't care as the MC locking is easy.

If we can help to reduce the arm motion with the MCL feedforward trained with an arm sometime before,
this means the lock acquisition will become easier. And this may still be compatible with the ALS.

Why did you notched out the 16Hz peak? It is the dominant component for the RMS and we want to eliminate it.

  7766   Fri Nov 30 11:38:35 2012 SteveUpdatePSLfoil removed from enclosure

Aluminum foil replaced by sheet metal on Enclosure and AP table.

Attachment 1: IMG_1822.JPG
IMG_1822.JPG
  7765   Fri Nov 30 09:59:53 2012 SteveHowToGeneralHow not to

Clean cabinet S15 doors were left open. You have to lock it up!

  7764   Fri Nov 30 02:40:44 2012 DenUpdateAdaptive FilteringYARM

I've applied FIR adaptive filter to YARM control. Feedback signal of the closed loop was used as adaptive filter error signal and OAF OUT -> IN transfer function I assumed to be flat because of the loop high gain at low frequencies. At 100 Hz deviation was 5 dB so I've ignored it.

I've added a filter bank YARM_OAF to C1LSC model to account for downsampling from 16 kHz to 2 kHz and put low-pass filter inside.

I've used GUR 1&2 XYZ channels as witnesses. Bandpass filters 0.4-10 Hz we applied to each of them. Error signal was filters using the same bandpass filter and 16 Hz 40 dB Q=10 notch filter. As an AI filter I used 32 Hz butterworth 4 order low-pass filter. Consequently, AI, bandpass and notch filters were added to adaptive path of witness signals.

I've used an FIR filter with 4000 taps, downsampling = 16, delay = 1, tau = 0, mu = 0.01 - 0.1. Convergence time was ~3 mins.

yarm_oaf.png

  7763   Thu Nov 29 09:58:06 2012 DenUpdatePEMDecreased RMS in Seismometers

Quote:

[Den, Ayaka]

We found that seismometer was working and the calibration in the filter banks should have been wrong.
We turned off the all FM2 filter in RMS filter banks.
 

We also installed STS seismometer. It is under the BS. Now we have spectrum of three seismometers;
 

RA: the above plot is kind of unreadable and useless. Please replace with something legible and put in some words about why there is a wrong filter, what exactly it is, etc., etc. etc. And why would you leave in a filter which is not supposed to be on? We might as well leave a few secretly broken chairs in the control room...

 First of all, STS-2 is in the end of X arm, GUR2 is under BS, GUR1 is in the end of Y arm.

BLRMS were small because we applied calibration from counts to um/s two times. In the past we had calibration in the RMS BP filter bank (vel2vel = FM2). Now we have calibration in the seismometer input filter bank so we can save calibrated _OUT channels.

  7762   Thu Nov 29 02:43:48 2012 AyakaUpdatePEMDecreased RMS in Seismometers

Quote:

The attached plots display RMS noise from various accelerometers and seismometers over the past 90 days. One can see how after the reinstallation of the seismometers in November, RMS from the GUR1Z and GUR1X channels decreases by a factor of about 100 from data in August. Additionally, the RMS over the course of the last 90 days has notably decreased in all instruments. In many cases, the RMS is only the result of inherent electronics noise, rather than from a signal.

The Image is replaced

[Den, Ayaka]

We found that seismometer was working and the calibration in the filter banks should have been wrong.
We turned off the all FM2 filter in RMS filter banks.

We also installed STS seismometer. It is under the BS. Now we have spectrum of three seismometers.
GUR1Xfilterbank.pngseismometers1129.pdf


 

RA: the above plot is kind of unreadable and useless. Please replace with something legible and put in some words about why there is a wrong filter, what exactly it is, etc., etc. etc. And why would you leave in a filter which is not supposed to be on? We might as well leave a few secretly broken chairs in the control room...

Attachment 2: seismometers1129.jpg
seismometers1129.jpg
  7761   Thu Nov 29 00:15:13 2012 Den, KojiUpdateIOOMC WFS work

Quote:

The instability started about 48hour ago, that means my work on the AP table did not 
made immediate trouble. But it does not mean anything.

For now, the WFS outputs are off. More work is needed to find what's wrong.

 

 The problem was caused by low reflectivity of the mirror that splits MC reflected beam into two: first goes to trash, second - to WFS. Power before the mirror was 100mW, reflected beam that goes to WFS was 0.3mW. Using dataviewer we learnt that the beam intensity was ~5 times more in the past.

This happened because the mirror position was adjusted a few days ago. Its reflection depends on the angle of incidence and amount of light to WFS was significantly reduced. We could either increase the angle of incidence or use two mirrors with high reflectivity instead of this with high transmission.

We've chosen the second variant not to confuse anyone in future with non-45 degrees angles. We are now using one mirror with reflectivity 98% to direct most power to the trash while other 2% are directed using the second mirror to WFS path. We now have 0.7 mW on WFS1 and 1.3 mW on WFS2.

Then we adjusted WFS 

  • blocked the beam and run scripts/MC/WFS/WFS_FilterBank_offsets to calculate offsets in the WFS servo
  • aligned MC and centered beams on WFS 1 and 2
  • provided excitation to MC1 at 5 Hz (400 counts) and adjusted I&Q phase rotation
  • adjusted the gain and changed it in MC autolocker (reduced from 0.25 to 0.15 as we now have more power of WFS as before)

We were able to close the loops. The phase margin is too low though, we need to improve feedback filters.

DSC_4942.JPG

Attachment 1: wfs_fb.pdf
wfs_fb.pdf wfs_fb.pdf wfs_fb.pdf wfs_fb.pdf wfs_fb.pdf wfs_fb.pdf
  7760   Wed Nov 28 23:55:13 2012 AyakaUpdateWienerFilteringSound on PSL

 I have been searching for the way we can subtract signal better since I could see the acoustic coupling signal remains in the target signal even though there are no coherence between them.

I changed the training time which is used to decide wiener filter.
I have total 10 minutes data, and the wiener filter was decided using the whole data before.
tri40_psdcoh_mcf_varioust.pnggauss_psdcoh_mcf_varioust.png

(Right: the performance with the data when the triangular sound was created. Left: the performance with the data when the gaussian sound was created.)

I found that the acoustic signal can be fully subtracted above 40 Hz when the training time is short. This means the transfer functions between the acoustic signals and MCF signal change.
However, if the wiener filter is decided with short-time training, the performances at lower frequencies get worse. This is because wiener filter do not have enough low-frequency information.

So, I would like to find the way to combine the short-time training merit and long-time training merit. It should be useful to subtract the broad-band coupling noise.

  7759   Wed Nov 28 23:18:35 2012 CharlesUpdatePEMDecreased RMS in Seismometers

The attached plots display RMS noise from various accelerometers and seismometers over the past 90 days. One can see how after the reinstallation of the seismometers in November, RMS from the GUR1Z and GUR1X channels decreases by a factor of about 100 from data in August. Additionally, the RMS over the course of the last 90 days has notably decreased in all instruments. In many cases, the RMS is only the result of inherent electronics noise, rather than from a signal.

Attachment 1: 8-31_11-29_PEM-RMS.jpg
8-31_11-29_PEM-RMS.jpg
  7758   Wed Nov 28 21:42:21 2012 ranaFrogsComputer Scripts / Programsdataviewer font error

An error this evening on rossa: dataviewer not working due to some font errors:

controls@rossa:~ 0$ dataviewer
Connecting.... done
Warning: Not all children have same parent in XtManageChildren
Warning: Not all children have same parent in XtManageChildren
Warning: Not all children have same parent in XtManageChildren
Warning: Not all children have same parent in XtManageChildren
Warning: Not all children have same parent in XtManageChildren
Warning:
    Name: FilterText
    Class: XmTextField
    Character '\52' not supported in font.  Discarded.

Warning:
    Name: FilterText
    Class: XmTextField
    Character '\56' not supported in font.  Discarded.

Warning:
    Name: FilterText
    Class: XmTextField
    Character '\170' not supported in font.  Discarded.

Warning:

etc.............

  7757   Wed Nov 28 17:40:28 2012 jamieOmnistructureComputerselog working again on firefox 17

Koji and I figured out what the problem is.  Apparently firefox 17.0 (specifically it's user-agent string) breaks fckeditor, which is the javascript toolbox the elog uses for the wysiwyg text editor.  See https://support.mozilla.org/en-US/questions/942438.

The suspect line was in elog/fckeditor/editor/js/fckeditorcode_gecko.js.  I hacked it up so that it stopped whatever crappy conditional user agent crap it was doing.  It seems to be working now.

Edit by Koji: In order to make this change work, I needed to clear the cache of firefox from Tool/Clear Recent History menu.

  7756   Tue Nov 27 19:06:16 2012 RijuUpdate Testing AG4395A+GPIB

 I ve tested another bandpass filter today with similar set-up. This time I took the data with corrected reference level. To set this reference-level the filter was disconnected and the cable was connected "thru" according to the instructions provided in the manual of AG4395A at http://cp.literature.agilent.com/litweb/pdf/04395-90040.pdf, page 3-10. The transfer functions are as follows 

Attachment 1: tfmag1.jpg
tfmag1.jpg
Attachment 2: tfphase1.jpg
tfphase1.jpg
  7755   Tue Nov 27 17:45:42 2012 SteveUpdateGeneralspare optics of AP table moved to cabinet

Spare optics from the AP table were moved to glass cabinet in the east arm. I'm not sure this is the right place. We'll see what everybody thinks.

There were two UNMARKED optics. Shame on you! No pencil marks on the optics either. These optics were shipped to the FBI for finger tip analysis.

 

 

Attachment 1: IMG_1819.JPG
IMG_1819.JPG
  7754   Tue Nov 27 11:27:22 2012 SteveUpdatePEMparticle count is going up

Jeff is still working on the filter banks.
Attachment 1: 5yparticles.png
5yparticles.png
Attachment 2: particle_6m_6m.png
particle_6m_6m.png
Attachment 3: 1y_particlec.png
1y_particlec.png
Attachment 4: 10y_particle_c.png
10y_particle_c.png
  7753   Tue Nov 27 09:44:09 2012 SteveUpdateVACpumpdown at 15 days

 It took 15 days to reach 1e-5 Torr

The RGA was turned on yesterday. The rga scan plotting was not working. Den wrote a Python script to log and plot because the Mathlab had trouble receiving data from our old c0raga.

It turned out that the Joe's log&plot system was working today. Thanks Den and Joe.

Here is the first rga scan.

 

Attachment 1: pd74m15d.png
pd74m15d.png
Attachment 2: pd74m15d_rga1d.png
pd74m15d_rga1d.png
  7752   Tue Nov 27 03:26:00 2012 JenneUpdateGeneralPOP beam on POP55, POP camera

Quote:

Some obvious things to be fixed

-  We need POP55 and POP CCD for diagnoses.

 Done.  The beam is also going vaguely in the direction of POP110, but I can't see the beam, so it's tricky. 

Order of operations:

1. Find POP on the table, place iris so I wouldn't forget.  Find beam by putting big IR card where I think beam should be, look at IR card with IR viewer.

2. Move and re-clamp 2" lens so beam is on center of lens.

3. Move and re-clamp 1st (2") mirror so that beam is on center of mirror. 

4. Remove BS-33% so that all the beam goes to POP55, steer that 1st mirror so beam is on POP55's little mirror.  Align little mirror so beam is centered on POP55 (as seen by looking at PD with viewer, finding "edges" of PD, going back to center).

5. Put BS-33% back in place.  The reflected portion of this beam is not possible to see using card+viewer technique.

6. Remove BS-50% that reflects half of this beam to POP110.  Find beam reflected from BS-33% by waving POP camera around.  Steer BS-33 until beam goes back in the direction that the camera used to be mounted.  Adjust camera mount and BS-33 so that beam is on camera.

7. Put BS-50% back in place.  Steer it around with voltmeter on PD to see if beam ever hits PD.  Unsuccessful. Give up, since we have POP55, and POP camera.

8. Make a youtube video: POP, AS, REFL, ITMXF (all on Quad3) - PRMI coarsely aligned, no IFO parts locked.  MICH was locked earlier, but not during video time.

  7751   Tue Nov 27 01:03:42 2012 AyakaUpdateWienerFilteringSound on PSL

 Last Thursday, I put the speaker and my laptop in the PSL table, and make triangular wave sound with the basic frequency of 40Hz, and Gaussian distributed sound.
(I create the sounds from my laptop using the software 'NHC Tone Generator' because I could not find the connector from BNC to speaker plug.)
And I measured the acoustic coupling in MCF signal. The all the 6 microphones were set in PSL table around PMC and PSL output optics. 

The performance of the offline noise cancellation with wiener filter is below.
(The target signal is MCF and the witness signals are 6 microphones.)

  • With Gaussian sound (Sorry for wrong labeling 'XARM' and no calibration)
    gauss_psdcoh_mcf.png
  • With 40Hz Triangular sound (Sorry for no calibration again)
    tri40_psdcoh_mcf.png

I can see some effects on MCF due to the sound on PSL table. Though I can subtract some acoustic signal and there are no coherence between MCF signal and mic signals, still some acoustic noise remains.
This is maybe because of some non-linearity effects or maybe because we have other effective places for acoustic coupling measurement. More investigations are needed.

Also, I compared the wiener filter and the transfer function from microphones signal to MCF signal. They should be the same ideally.
gauss_filters_mcf.pngGauss_estimatefilters_mcf.png

(Left: Wiener filter, Right: Transfer function estimated by the spectrum. They are measured when the Gaussian sound is on.)

These are different especially lower frequencies than 50 Hz. The wiener filter is bigger at lower frequencies. I guess this adds extra noise on the MCF signal. (see the 1st figure.)
The wiener filter can be improved by filterings. But if so, I want to know how can we determine the filters. It is interesting if we have some algorithms to determine the filters and taps and so on.
The more investigations are also needed.

  7750   Tue Nov 27 00:45:20 2012 jamieUpdateIOOMC_L and laser frequency noise spectra

I grabbed the a plot of the iLIGO PSL frequency noise spectrum from the Rana manifesto:

laser_noise.pdf

Rana's contention is that this spectrum (red trace) is roughly the same as for our NPRO.

From the jenne/mevans/pepper/rana paper Active noise cancellation in a suspended interferometer I pulled a plot of the calibrated MC_L noise spectrum:

MCL_noise.pdf

The green line on this plot is a rough estimate of where the above laser frequency noise would fall on this plot.  The conversion is:

    L / f  =  10 m / 2.8e14 Hz = 3.5e-14 m/Hz

which at 10 Hz is roughly 1.5e-11 m.  This puts the crossover somewhere between 1 and 10 Hz.

  7749   Tue Nov 27 00:26:00 2012 jamieOmnistructureComputersUbuntu update seems to have broken html input to elog on firefox

 After some system updates this evening, firefox can no longer handle the html input encoding for the elog.  I'm not sure what happened.  You can still use the "ELCode" or "plain" input encodings, but "HTML" won't work.  The problem seems to be firefox 17.  ottavia and rosalba were upgraded, while rossa and pianosa have not yet been.

I've installed chromium-browser (debranded chrome) on all the machines as a backup.  Hopefully the problem will clear itself up with the next update.  In the mean time I'll try to figure out what happened.

To use chromium: Appliations -> Internet -> Chromium

  7748   Mon Nov 26 23:45:52 2012 AyakaUpdateIOOTuning MC_L

[Rana, Jamie, Ayaka]

We could not lock the arms with MC_L loop on. Therefore we measured the change in YARM error signal when the MC_L is turned on.

POYerr_MCL.jpg

(data; POYerr_MCF.xml in zip file)

Green line; POY error signal when MCL loop was on and the YARM loop gain (0.5) was so high that the saturated control signal made funny peak around 250 Hz.
Blue line; POY error signal when MCL loop was off and the YARM loop gain was low (0.2).
Pink line; POY error signal when MCL loop was on (the gain was -300) and the YARM loop gain was low (0.2).
Red line; POY error signal when MCL loop was on, another low pass filter (2nd order, cut off frequency of 55Hz) was added to MCL loop and the YARM loop gain was low (0.2).

We also changed the filter trigger in order to lock YARM. The FM7 and 8 trigger was turned off. It means that spectrum above was took with FM2,3,4,5,6,9,10 on. Whitening filters were also on.

MCL control signal makes the arm spectrum bad because the MCL control signal moves MC2 mirror additionally and adds extra frequency noise.
Ideally, error signal should be the same at higher frequencies and go down at the lower frequencies when the MCL loop is on because MCL signal should suppress the seismic noise.

Before we added the LPF, MCF/MCL loop cross over (which was taken with the template /users/Templates/MC/MCL-MCF_xover-2012-8-23.xml) is below;

MCL-MCFxover.jpg

(MCL-MCF_xover.xml in zip file)

After the LPF is added, the cross over has been changed as below;
MCL-MCFxover2.jpg

(MCL-MCF_xover2.xml in zip file)

For now, I will just turn off the MCL loop for the acoustic noise experiments.

Attachment 4: 121126.tar.gz
  7747   Mon Nov 26 19:27:59 2012 RijuHowTo Testing AG4395A+GPIB

Riju, Jenne

We have checked the transfer function of a bandpass filter using AG4395A network analyzer and retrieved the data through GPIB. The RF out signal of AG4395A had been divided by splitter with two outputs of the splitter going to through R and the filter which was connected to the A channel of the network analyzer. The GPIB data came in complex data format, from which the absolute value and phase had to be retrieved. 

 

The plot for the TF is as following

Attachment 1: tfmag.jpg
tfmag.jpg
Attachment 2: tfphase.jpg
tfphase.jpg
  7746   Mon Nov 26 18:56:34 2012 JenneHowToComputersData logging suggestions

We've been talking for a while about how we want to store data.  I'm not in love with keeping it on the elog, although I think we should always be able to reference and go back and forth between the elogs and the data.

I have made a new folder: /data    EDIT: nevermind.  I want it to be on the file system just like /users, but I don't know how to do that.  Right now the folder is just on Ottavia. Jamie will help me tomorrow.

In this folder, we will save all of the data which goes into the elog. 

I propose that we should have a common format for the names of the data files, so that we can easily find things.

My proposal is that one begins ones elog regarding the data to be saved, and submit it immediately after putting in the first ~sentence or so. One should then make a new folder inside the data folder with a title "elog#####_Anything_Else_You_Want" Then, data (which was originally saved in ones own users folder) should be copied into the /data/elog#####_AnythingElse/ folder. Also in that folder should be any Matlab scripts used to create the plots that you post in the elog.  One should then edit the elog to continue making a regular, very thorough elog, including the path to the data.  Elog should include all of the information about the measurement, state of the IFO (or whatever you were measuring), etc. 

Riju will be alpha-testing this procedure tonight.  EDIT: nevermind...see previous edit.

  7745   Mon Nov 26 18:36:17 2012 JenneUpdatePEMBLRMS back

Quote:

 I got two seismometers and one microphone back from Tara.

They are now near the Gurlap under the MC.

 I have finally plugged GUR1 back in....it is down at ETMY for now, since that's where the cable was.  BLRMS are back up on the projector.

  7744   Mon Nov 26 10:58:20 2012 SteveUpdatePEMair cond maitenance tomorrow

Air conditioning maintenance is scheduled for tomorrow morning till noon.

  7743   Mon Nov 26 10:42:06 2012 AyakaUpdateLSCcalibration of arms

I uploaded a zip file that contains data files used for the calibration.

OLTF_x/y.txt: the open loop transfer function (measured by IN1/IN2 in arm servo filter bank).
coh_x/y.txt: coherence of OLTF. I used the data where coherence > 0.98.
ext_err_x/y.txt: the transfer function from ITM excitation signal to POX/POY error signal.
coh_x2/y2.txt: coherence of ext_err. I used the data where coherence > 0.98.

The LSC whitening filter was off because the xarm was unlocked when the POX Q whitening filter was turned on. (We have to study what was wrong.)
The SUS whitening filters were on.
The all digital filters except +6dB filter were on.

Attachment 1: armcalib.zip
  7742   Mon Nov 26 10:06:51 2012 ranaSummaryIOOMC slides from 2002
Attachment 1: MCtalk.pdf
MCtalk.pdf MCtalk.pdf MCtalk.pdf MCtalk.pdf MCtalk.pdf MCtalk.pdf
  7741   Sat Nov 24 23:50:35 2012 KojiUpdateIOOMC WFS refusing to work

Today I found the IMC was misaligned significantly by WFS feedback.
Once the feedback was cleared, it locks with nice visibility.
But WFS misaligns it again as soon as the intergrators are engaged.
I checked the beam on the table, but found nothing really wrong.
The offsets of the error signals were nulled at the input filter modules of the WFS segments.
They did not fix the problem.

The instability started about 48hour ago, that means my work on the AP table did not 
made immediate trouble. But it does not mean anything.

For now, the WFS outputs are off. More work is needed to find what's wrong.

 

  7740   Sat Nov 24 22:14:08 2012 KojiUpdateGeneralAP table cleaning up

On Wednesday (21st) night, I checked the AP table as I wanted to try PRMI locking. 
It was difficult to work with the table as there were so many unnecessary components on it.
Also the beams went through complicated paths as they have funny angles. 

So I decided to clean up of IMC REFl WFS, IFO REFL, and IFO AS paths.
I found that the AS beam was highly astigmatic as the beam went through a (too-much-) tilted lens.
I made several blocked optical paths for REFL and AS for future extension of the detection system.

The current status of the table was uploaded below.

The optical spectrum analyzers and the aux NPRO were left untouched but they should be moved
somewhere (either on the table or outside) which does not disturb the other optical paths.
 


After the cleaning, I started locking PRMI. I could lock PRMI stably. But I could not figure out how
the intra-cavity mode looked like as I did not have the POP camera. The power recycling
gain was not quantitatively evaluated as I did not have POP and I wasn't sure how the beam was aligned at POX/POY.


We need to know:

- Quantitative evaluation of the beam shape in the PRC

- Quantitative evaluation of the power recycling gain

Some obvious things to be fixed

- The POX whitening filters seem not switching. This issue should be checked at the circuit module itself and at the BIO.

- The POX beam is not well focused on the PD. This was particularly clear when PRMI was locked with carrier.

- The POP beam is going nowhere. We need POP55 and POP CCD for diagnoses.

I haven't checked ITMY table.

Attachment 1: APTable121124.pdf
APTable121124.pdf
Attachment 2: APTable121124.graffle.zip
  7739   Sat Nov 24 13:58:07 2012 ranaUpdateLSCcalibration of arms

For the loop diagnosis, its best to use the method of "IN1/IN2", rather than manipulate the close loop gain. In this way, you can directly plot the swept sine measurement from DTT as the open loop gain.

Also, for reporting calibration, we should all try to record the current settings better. Anything that may change the loop gain should be recorded along with the Bode plot and the DATA must be posted to the elog - no more of just posting plots.

We need to know, e.g.

  1. what is the power in the arms?
  2. are the LSC whitening filters on?
  3. are the SUS dewhitening filters on?
  4. What normalization is being used in the LSC?
  5. What digital filters are on in the X/YARM loop filter bank?

Resistance is feudal.

  7738   Wed Nov 21 21:06:13 2012 AyakaUpdateLSCcalibration of arms

Motivation

In order to estimate whether we can see acoustic coupling in arms or not, we have to calibrate signals to phase noise.

Method

I used the same method as Yuta and Jenne did (6834).
I switched from ETM locking to ITM locking since only ITM actuators are calibrated (5583), and measured the open loop transfer function and the transfer function from ITM excitation to POX/POY error signal. Then I can estimate the calibration value H [counts/m] from POY/POX error signal to displacement.

Results

Yarm; H = 9.51 x 1011 counts/m
  OL_y.pngerr_exc_y.pngPOY_disp_y.png

Xarm; H = 6.68 x 1011 counts/m
OL_x.pngerr_exc_x.pngPOX_disp_x.png

Phase noise in arms:
XY_phase.png
blue; Xarm, green; Yarm

 

Next Step

I will calibrate the acoustic signal and see if it is reasonable that we can see the acoustic coupling signal in the arms.
But I guess it is difficult. Actually I have not seen coherence between ETM feedback signals and acoustic sounds yet. (I measured acoustic noise near POX and in PSL table.)

If I find that it is hopeless, I will create some sounds and try to measure transfer function from acoustic sound to arm cavity signals.
I am interested in how the transfer function calculated by wiener filtering is different from the measured transfer function.

 

Note

I found that we do not have enough phase margin. This is why the arm locking is not so stable.

  7737   Wed Nov 21 01:36:36 2012 KojiUpdateIOOMCL disabled / WFS clear history restored

As MCL is disturbing arm locking by injecting a lot of noise, I have modified 'mcup'  to disable MCL

As MC WFS keeps failing to start up when it is locked, the lines in 'mcwfsoff' to clear WFS filter history were restored.

ELOG V3.1.3-